I have the problem:
41-8x= -6x-23
They say that x= -9.
Is this true or false?

Answer by mananth(12270)

41-8x= -6x-23
plugvalue of x=-9
41-8*(-9)=-6*(-9)-23
41+72= 54-23

113=31
so x is not equal to -9
...
m.ananth@hotmail.ca